OpenStudy (anonymous):

Are the following geometric sequences? 3, 9, 27, 81,… 1.5, 2.25, 3.375, 5.0625,…

OpenStudy (anonymous):

geometric means you can either multiply or divide each thing by the same number. like 3*3=9 9*3= 27 27*3=81 so that one is a yes! :)

OpenStudy (anonymous):

5.0625/1.5=3.375 3.375/1.5=2.25 and 2.25/1.5=1.5 :) What school do you go to, b/c I just had the same problems I did them last night lol

OpenStudy (anonymous):

I go to Falcon Virtual Academy. And thanks. :)

OpenStudy (anonymous):

oh yeah not my school lol.. But that's so weird! An dyou're welcome! :D